# On the Origin of the High Star-Formation Efficiency in Massive Galaxies at Cosmic Dawn

### v1

@@ 2.2 Subgrid turbulence model
We use a Large Eddy Simulation (LES) method for subgrid turbulence similar to Kretschmer & Teyssier (2020). To describe this model, we decompose the subgrid density field $\rho$ into a bulk component $\overline{\rho}$ and a turbulent component $\rho^{\prime}$:
@@
The bulk components are defined as the mass-weighted volume-average fields, also known as Favre-averaged fields.
The specific TKE is defined as
@@
The bulk fluid equations include new terms for turbulent diffusion. However, the turbulent diffusion is generally small compared to the diffusion introduced by the numerical scheme. We maintain the original fluid equations rather than adding a diffusive term to an already too-diffusive scheme, following previous work (Schmidt et al., 2006; Kretschmer & Teyssier, 2020).
The TKE equation reads
@@
where $P_{\rm turb}=2/3\varepsilon_{\rm turb}=\overline{\rho}\sigma_{\rm 1D}^{2}$ is the turbulent pressure and $\mathcal{C}_{\rm turb}$ and $\mathcal{D}_{\rm turb}$ are creation and destruction source terms respectively.
@@ 2.3.1 Subgrid density PDF
Empirically, star formation is described by the Kennicutt-Schmidt (KS) law (Kennicutt, 1998), which relates the SFR surface density and gas surface density by $\Sigma_{\rm SFR}\propto\Sigma_{\rm gas}^{1.5}$. The KS relation holds at scales as small as $1~{}{\rm kpc}$ (Kennicutt et al., 2007), supporting the idea that star formation can be modeled locally as volumetric Schmidt law
@@
$p(s)=\frac{1}{\sqrt{2\pi\sigma_{s}^{2}}}\exp(-\frac{(s-\overline{s})^{2}}{2\sigma_{s}^{2}})$ (10)
@@ 2.3.2 Turbulence forcing parameter
In most of our models, we set the forcing parameter to a constant value. However, we run one experimental model where we derive the forcing parameter from the local velocity field. Using a Helmholtz decomposition, the velocity field can be represented as a curl-free term $\mathbf{v}_{\rm comp}=-\gradient{\phi}$ and divergence-free term $\mathbf{v}_{\rm sol}=\curl{\mathbf{A}}$.
@@
Federrath et al. (2010) fit a simple analytic form to $b_{\rm turb}$ using detailed turbulence simulations (their Eq. 23):
@@ 2.3.3 Star formation efficiency
Assuming that star-forming cores are homogeneous spheres with diameter $\ell_{\rm s}$ (Krumholz & McKee, 2005), the gravitational stability condition is $\alpha_{\rm vir,cl}\geq 1$, where the virial parameter $\alpha_{\rm vir,s}$ is
$\alpha_{\rm vir,cl}=\frac{2E_{\rm kin}}{\absolutevalue{E_{\rm grav}}}=\frac{15}{\pi}\frac{c_{\rm s}^{2}+\sigma(\ell_{\rm s})^{2}}{G\rho\ell_{\rm s}^{2}}$ (16)
The gravitational stability condition can alternatively be expressed as a condition on the density $\rho\leq rho_{\rm crit}$, where
@@
where
$\alpha_{\rm vir}=\frac{15}{\pi}\frac{c_{\rm s}^{2}+\sigma_{\rm 1D}^{2}}{G\overline{\rho}\Delta x^{2}}=\frac{15}{\pi}\frac{c_{\rm s}^{2}}{G\overline{\rho}\Delta x^{2}}(1+\mathcal{M}^{2})$ (18)
@@
$s_{\rm crit}=\ln\left[\alpha_{\rm vir}\frac{2\mathcal{M}_{\rm turb}^{4}}{1+\mathcal{M}^{2}}\right]$ (19)
By design, $s_{\rm crit}$ depends on the cell size. At higher resolutions, a larger density is required to become gravitationally unstable, but gas will naturally reach those higher densities as it collapses down to that smaller cell size.
The model breaks down when the entire cell is subsonic $\mathcal{M}_{\rm turb}\lesssim 1$ (Federrath & Klessen, 2012) and the sonic scale is larger than the resolution scale. In this case, the density PDF should be interpreted as the probability distribution for the density across the entire cell and the gravitational stability condition (Equation 16) should use the resolution scale rather than the sonic scale. We smoothly interpolate between both stability conditions by defining a modified critical logarithmic density (Kretschmer & Teyssier, 2020):
@@
If each gas parcel that does not meet the stability condition collapses in one freefall time and converts all its mass into stars, then the local SFR is given by
@@
For notational convenience, we drop the diacritical marks on the bulk fluid components for the remainder of the paper.
This is the multi-freefall (MFF) model of Federrath & Klessen (2012). This recipe is an important first step towards predictive models at Cosmic Dawn. However, there are several caveats. The model neglects deviations from a lognormal density PDF, the spatial correlation of the density field, and the timescale required to replenish the density PDF after a star has formed. Some of these issues are addressed by the recent turbulent support (TS) model of Hennebelle et al. (2024) (Sec. 4.6).
The MFF model implies that as the turbulence forcing becomes more solenoidal, the subgrid density PDF becomes more narrow (Eq. 12), generally reducing the local SFE for the same gas conditions and TKE (Eq. 22). We leave further discussion to Section 4.4.
The expected amount of star cluster particles to form in one timestep is
@@
where $M_{\rm cl}$ is the star cluster particle mass. In our simulations, the fiducial value $M_{\rm cl}=M_{\rm b,min}=4400~{}M_{\odot}$ is the minimum baryon mass resolution on the coarse grid defined in Section 2.1. We draw the number of star cluster particles to form from a Poisson distribution with a Poisson parameter $\expectationvalue{N_{\rm cl}}$. This procedure allows us to model star formation as discrete events distributed in time, independent of the adopted star cluster particle mass.
@@ 2.4 Supernovae recipe
SN explosions are one of the key feedback processes which regulate star formation. As the SFR increases, the SN rate increases commensurately. SNe inject energy and momentum into the ISM, depleting the reservoir of cold gas available to form stars.
@@
Each star cluster particle is expected to produce a number of SNe given by
@@
where $M_{\rm cl}$ is the mass of the star cluster particle, $m_{\rm big}$ is the typical mass of a massive star, and $\chi$ is the mass fraction of massive stars within the stellar population. $m_{\rm big}$ and $\chi$ are related to the IMF $\xi(m)$ via
@@
Our choice of $m_{\rm big}$ and $\chi$ are the only way in which the IMF enters into our modeling. Otherwise, our model implicitly assumes that our star cluster particles fully sample the IMF. Using a Chabrier IMF (Chabrier, 2003), one finds $m_{\rm big}\simeq 19~{}M_{\odot}$ and $\chi\simeq 0.19$. We adopt $m_{\rm big}=10~{}M_{\odot}$ and $\chi=0.2$, producing somewhat more SNe than would be predicted by the Chabrier IMF, but only by a small factor well within the uncertainties of the IMF at Cosmic Dawn.
@@
Multiplying by the fine timestep $\Delta t$, we obtain the expected number of SN events per timestep $\expectationvalue{N_{\rm SNe}}=\dot{N}_{\rm SNe}\Delta t$. We draw the number of SN events from a Poisson distribution with a Poisson parameter $\lambda=\expectationvalue{N_{\rm SNe}}$, similar to Hopkins et al. (2018). This procedure allows us to model SNe as discrete events distributed in time, independent of the adopted star cluster particle mass.
The an exploding SN blastwave interacts with the surrounding medium in 4 stages. Initially, the SN ejecta expands freely. Once the ejecta has swept up a mass in the ISM comparable to its own, a shock forms and expands following the Sedov-Taylor solution, approximately conserving energy. Eventually, radiative cooling behind the shock front causes the expansion to deviate from the Sedov-Taylor solution, and the expansion proceeds approximately conserving momentum. Finally, the blast wave slows due to radiative losses and accumulated material, fading into the ISM. The transition radius between the second Sedov-Taylor phase and the third “snowplow” phase is called the cooling radius.
Ideally, one should resolve the energy-conserving Sedov-Taylor phase and model a SN explosion by injecting a thermal energy $E_{\rm SN}\simeq 10^{51}~{}{\rm erg}$ into the surrounding gas. However, at high densities, the cooling radius is often smaller than the resolution scale. In this case, injecting energy into the surrounding gas would result in the energy from the SN being artificially radiated away before it could accelerate gas in the snowplow phase. To compensate for this effect, we inject momentum in addition to thermal energy when the cooling radius is not resolved. This method, known as momentum feedback, is used frequently in galaxy formation simulations (Hopkins et al., 2011, 2018; Kim & Ostriker, 2015; Kimm et al., 2015).
@@
$R_{\rm cool}=$ $3.0~{}{\rm pc}\left(\frac{Z}{Z_{\odot}}\right)^{-0.082}\left(\frac{n_{\rm H}}{100~{}{\rm cm^{-2}}}\right)^{-0.42}$ (27)
$P_{\rm SN}=$ $1.42\times 10^{5}~{}{\rm km/s}M_{\odot}\left(\frac{Z}{Z_{\odot}}\right)^{-0.137}\left(\frac{n_{\rm H}}{100~{}{\rm cm^{-2}}}\right)^{-0.16}$ (28)
@@
Our approach is only an approximation of the full model recommended by Martizzi et al. (2015), which incorporates additional physical scales into the calculation of injected momentum and energy. Martizzi et al. (2015) also provide versions of Eqs. 27 and 28 for SN which occur in an inhomogeneous medium with $\mathcal{M}_{\rm turb}=30$. In an inhomogeneous medium, the cooling radius is larger by a factor of a few because the blast wave preferentially moves through low-density channels, but the terminal energy and momentum deposited in the ISM by the SN remnant do not change significantly. Ideally, one should use a model which depends continuously on the turbulent Mach number, no such model is available at the time of writing.
@@
Although momentum feedback approximates the net effect of SNe on the surrounding ISM, it does not accurately model the effect of SNe on the surrounding ISM on short timescales. Therefore, our results concerning SN feedback should be considered with caution, despite our physically-motivated subgrid prescription (Sec. 4.7).
@@ 2.5 Early feedback recipe
In addition to SN feedback, stellar feedback includes thermal pressure from photoionized gas, stellar winds, radiation pressure from FUV photons, and radiation pressure from multiply-scattered IR photons reprocessed by dust. We call these early feedback processes because they begin immediately after a massive star forms, in contrast to SN feedback. Out of these feedback processes, we only model thermal pressure from photoionized gas. We discuss the effects of other early feedback processes in Section 4.5.
In each cell containing a star cluster particle younger than $\tau_{\rm end}=20~{}{\rm Myr}$, we set the temperature to $T_{\rm phot}=20000~{}{\rm K}$. This forces the gas into a fast cooling regime where it approaches the photoionization temperature within a few timesteps. In practice, we achieve this by injecting a thermal energy associated with a sound speed
@@
where we have assumed a $\gamma$ = 5/3 adiabatic equation of state and mean molecular weight $\mu=0.6$, appropriate for an ionized monoatomic gas.
Unlike our SN recipe, our early feedback recipe depends on the choice of $M_{\rm cl}$ because each star cluster particle has the same effect on its parent cell, regardless of its mass. If the adopted $M_{\rm cl}$ is small, then the stellar mass will be represented by a large number of star cluster particles which each photoionize their parent cell, representing efficient photoionization feedback. If the adopted $M_{\rm cl}$ is large, then the stellar mass will be represented by a small number of star cluster particles, representing inefficient photoionization feedback. $M_{\rm cl}$ should be chosen according to the stellar mass required to photoionize a cell. In this case, a star cluster particle is formed exactly when its HII region becomes resolved.
@@
Using data for main sequence stars given by Table 15.1 of Draine (2011), we fit $Q(m)$ to a simple analytic form for masses $M\geq 15.6~{}M_{\odot}$ (App. B):
$Q(m)\sim\left[8.37\times 10^{45}\left(\frac{m}{M_{\odot}}\right)^{2.12}-2.33\times 10^{48}\right]~{}{\rm cm^{-2}s^{-1}}$ (32)
Assuming a Chabrier IMF (Chabrier, 2003), we estimate $\expectationvalue{Q/m_{*}}=5.85\times 10^{46}~{}{\rm cm^{-2}s^{-1}M_{\odot}^{-1}}$. Higher mass stars contribute more ionizing photons, but have a lower abundance. In our calculation, most of the contribution to the ionizing photons comes from stars $M_{*}\simeq 35~{}M_{\odot}$. Star clusters which are too small to sample the IMF up to this mass may have lower ionization rates.
@@
where $\alpha_{B}\simeq 2.58\times 10^{-13}~{}{\rm cm^{3}/s}$ is the case B recombination rate at an electron temperature $T_{\rm e}=10^{4}~{}{\rm K}$ (Ferland et al., 1992). We use the on-the-spot approximation, assuming that electrons which recombine directly to the ground state do not contribute to the net ionization.
The stellar mass required to photoionize a cell is given by equating Strömgren sphere and cell volumes and solving for the star cluster mass, which yields
$\begin{split}M_{\rm cl}=\ &\frac{4\pi}{3}\frac{\alpha_{B}n_{\rm H}^{2}\Delta x_{\rm min}^{3}}{\expectationvalue{Q/m_{*}}}\\ \simeq\ &5000~{}M_{\odot}\left(\frac{n_{\rm H}}{100~{}{\rm cm^{-2}}}\right)^{2}\left(\frac{\Delta x}{10~{}{\rm pc}}\right)^{3}\left(\frac{\expectationvalue{Q/m_{*}}~{}{\rm cm^{2}sM_{\odot}}}{5.85\times 10^{-6}}\right)^{-1}\end{split}$ (34)
In MDGs, stars may form at significantly higher densities than the value $n_{\rm H}\sim 10^{2}~{}{\rm cm^{-2}}$ which appears in Equation 34. However, the strong turbulence in MDGs lowers the effective density relevant for photoionization because photoionizing photons preferentially escape through low-density channels in the gas. We can estimate the effective density using a Rosseland-like averaging
@@
A similar concept of effective density was used in Appendix B of (Faucher-Giguère et al., 2013). For a mean density $\sim 10^{4}~{}{\rm cm^{-2}}$ and a turbulent Mach number $\sim 10$, the effective density is indeed on the order of $\sim 10^{2}~{}{\rm cm^{-2}}$.
Equation 34 gives an order-of-magnitude estimate for an appropriate star cluster particle mass in our simulation. However, it depends sensitively on density, which our simple model cannot account for. Due to the steep density dependence and the uncertainty in $\expectationvalue{Q/m_{*}}$, we vary $M_{\rm cl}$ by over an order of magnitude in different simulations (Sec. 2.6). Although crude, this model provides a rough approximation for the effect of early feedback, and will be improved upon in future work.
One caveat of this model is that we do not properly account for multiple star cluster particles in a cell. If one star cluster particle photoionizes a cell, then by the Strömgren sphere argument, multiple star particles should photoionize a region larger than a cell. In our simulations, these situations are instead handled by multiplying the injected thermal energy by the number of star cluster particles, artificially restricting the Strömgren sphere to the cell size.

### v2

@@ 2.2 Subgrid turbulence model
We use a Large Eddy Simulation (LES) method for subgrid turbulence as implemented in RAMSES by Kretschmer & Teyssier (2020), which we describe here for clarity. To describe the model, we decompose the density field into a bulk component $\overline{\rho}$ and a turbulent component $\rho^{\prime}$:
@@
The bulk components are defined as the mass-weighted volume-average fields, also known as Favre-averaged fields. The specific TKE is defined as
@@
The bulk fluid equations include new terms for turbulent diffusion. However, the turbulent diffusion is generally small compared to the diffusion introduced by the numerical scheme. We maintain the original fluid equations rather than adding a diffusive term to an already too-diffusive scheme, following previous work (Schmidt et al., 2006). The TKE equation reads
@@
where $P_{\rm turb}=(2/3)\varepsilon_{\rm turb}=\overline{\rho}\sigma_{\rm 1D}^{2}$ is the turbulent pressure and $\mathcal{C}_{\rm turb}$ and $\mathcal{D}_{\rm turb}$ are creation and destruction source terms respectively.
@@ 2.3 Star formation recipe
With slight modification (Sec. 2.3.2), we model star formation using the multi-freefall model of Federrath & Klessen (2012) as implemented in RAMSES by Kretschmer & Teyssier (2020), which we describe here for clarity.
@@ 2.3.1 Subgrid density PDF
Empirically, star formation is described by the Kennicutt-Schmidt (KS) law (Kennicutt, 1998), which relates the SFR surface density and gas surface density by $\Sigma_{\rm SFR}\propto\Sigma_{\rm gas}^{1.5}$. The KS relation holds at scales as small as $1~{}{\rm kpc}$ (Kennicutt et al., 2007), supporting the idea that star formation can be modeled locally as volumetric Schmidt law (Schmidt, 1959)
@@
[…]
@@ 2.3.2 Turbulence forcing parameter
In most of our simulations, we set the forcing parameter to a constant value. However, we run one experimental simulation where we derive the forcing parameter from the local velocity field, a novel addition to RAMSES in this work. Using a Helmholtz decomposition, the velocity field can be represented as a curl-free term $\mathbf{v}_{\rm comp}=-\gradient{\phi}$ and divergence-free term $\mathbf{v}_{\rm sol}=\curl{\mathbf{A}}$.
@@
See Ginzburg et al. (2025) for a more rigorous justification of this expression. Federrath et al. (2010) fit a simple analytic form to $b_{\rm turb}$ using detailed turbulence simulations (their Eq. 23):
@@
We use this formula to compute a local turbulence forcing for each cell in the simulation.
@@ 2.3.3 Star formation efficiency
Assuming that star-forming cores are homogeneous spheres with diameter $\ell_{\rm s}$ (Krumholz & McKee, 2005), the gravitational stability condition is $\alpha_{\rm vir,core}\geq 1$, where $\alpha_{\rm vir,core}$ is the virial parameter of the core given by
$\alpha_{\rm vir,core}=\frac{2E_{\rm kin}}{\absolutevalue{E_{\rm grav}}}=\frac{15}{\pi}\frac{c_{\rm s}^{2}+\sigma(\ell_{\rm s})^{2}}{G\rho\ell_{\rm s}^{2}}$ (16)
The gravitational stability condition can alternatively be expressed as a condition on the density $\rho\leq\rho_{\rm crit}$, where
@@
and
$\alpha_{\rm vir}=\frac{15}{\pi}\frac{c_{\rm s}^{2}+\sigma_{\rm 1D}^{2}}{G\overline{\rho}\Delta x^{2}}=\frac{15}{\pi}\frac{c_{\rm s}^{2}}{G\overline{\rho}\Delta x^{2}}(1+\mathcal{M}_{\rm turb}^{2})$ (18)
@@
$s_{\rm crit}=\ln\left[\alpha_{\rm vir}\frac{2\mathcal{M}_{\rm turb}^{4}}{1+\mathcal{M}_{\rm turb}^{2}}\right]$ (19)
$s_{\rm crit}$ depends on the cell size. At higher resolutions, a larger density is required to become gravitationally unstable, but gas will naturally reach those higher densities as it collapses down to the smaller cell size.
The model breaks down when the entire cell is subsonic $\mathcal{M}_{\rm turb}\lesssim 1$ (Federrath & Klessen, 2012) and the sonic scale is larger than the resolution scale. In this case, the density PDF should be interpreted as the probability distribution for the density across the entire cell and the gravitational stability condition (Eq. 16) should use the resolution scale rather than the sonic scale. We smoothly interpolate between both stability conditions by defining a modified critical logarithmic density (Kretschmer & Teyssier, 2020):
@@
If each unstable gas parcel collapses in one freefall time and converts all its mass into stars, then the local SFR is given by
@@
For notational convenience, we drop the overlines on the bulk fluid components for the remainder of the paper.
This is the multi-freefall (MFF) model of Federrath & Klessen (2012), validated by detailed turbulence simulations in the same work. This recipe is an important first step towards predictive models at Cosmic Dawn. However, there are several caveats. The model neglects deviations from a lognormal density PDF, the spatial correlation of the density field, and the timescale required to replenish the density PDF after a star has formed. Some of these issues are addressed by the recent turbulent support (TS) model of Hennebelle et al. (2024) (Sec. 4.6). The MFF model implies that as the turbulence forcing becomes more solenoidal, the subgrid density PDF becomes more narrow (Eq. 12), generally reducing the local SFE for the same gas conditions and TKE (Eq. 22). We leave further discussion to Section 4.4.
The expected number of star cluster particles to form in one timestep $\Delta t$ is
@@
where $M_{\rm cl}$ is the star cluster particle mass. In our simulations, the fiducial value $M_{\rm cl}=M_{\rm b,min}=4400~{}M_{\odot}$ is the minimum baryon mass resolution on the coarse grid defined in Section 2.1. Each timestep, we form a number of star cluster particles sampled from a Poisson distribution with a Poisson parameter $\expectationvalue{N_{\rm cl}}$. This procedure allows us to model star formation as discrete events distributed in time, independent of the adopted star cluster particle mass.
@@ 2.4 Supernovae recipe
We use a SN feedback model based on the detailed turbulence simulations of Martizzi et al. (2015) as implemented in RAMSES by Kretschmer & Teyssier (2020), which we describe here for clarity. SN explosions are one of the key feedback processes which regulate star formation. As the SFR increases, the SN rate increases commensurately. SNe inject energy and momentum into the ISM, depleting the reservoir of cold gas available to form stars.
@@
Assuming that it samples the IMF, a star cluster particle is expected to produce a number of SNe given by
@@
where $M_{\rm cl}$ is the star cluster particle mass, $m_{\rm big}$ is the typical mass of a supernova progenitor, and $\chi$ is the mass fraction of supernova progenitors within the stellar population. $m_{\rm big}$ and $\chi$ are related to the IMF $\xi(m)$ via
@@
where $m_{\rm min}$ is the minimum mass of a SN progenitor and $m_{\rm max}$ is the maximum stellar mass.
The IMF only enters into our modeling through our choice of $m_{\rm big}$ and $\chi$. Using a Chabrier IMF (Chabrier, 2003) with $m_{\rm min}=8~{}M_{\odot}$ and $m_{\rm max}=100~{}M_{\odot}$, one finds $m_{\rm big}\simeq 19~{}M_{\odot}$ and $\chi\simeq 0.21$. Following Kretschmer & Teyssier (2020), we adopt $m_{\rm big}=10~{}M_{\odot}$ and $\chi=0.2$. Our smaller value of $m_{\rm big}$ results in approximately twice as many SNe than predicted by a Chabrier IMF. This discrepancy was unintentional. However, we note that the value of $m_{\rm big}$ is uncertain due to the sensitivity of Equation 25 to the exact choices of $m_{\rm min}$, $m_{\rm max}$, and the functional form of the IMF.
@@
The expected number of SN events per timestep is $\expectationvalue{N_{\rm SNe}}=\dot{N}_{\rm SNe}\Delta t$. Each timestep, we trigger a number of SNe sampled from a Poisson distribution with a Poisson parameter $\expectationvalue{N_{\rm SNe}}$, similar to Hopkins et al. (2018). This procedure allows us to model SNe as discrete events distributed in time, independent of the adopted star cluster particle mass.
The exploding SN blastwave interacts with the surrounding medium in 4 stages. Initially, the SN ejecta expands freely. Once the ejecta has swept up a mass in the ISM comparable to its own, a shock forms and expands following the Sedov-Taylor solution, approximately conserving energy. Eventually, radiative cooling behind the shock front causes the expansion to deviate from the Sedov-Taylor solution, and the expansion proceeds approximately conserving momentum. Finally, the blast wave slows due to radiative losses and accumulated material, fading into the ISM. The transition radius between the second Sedov-Taylor phase and the third “snowplow” phase is called the cooling radius.
Ideally, one should resolve the energy-conserving Sedov-Taylor phase and model a SN explosion by injecting a thermal energy $E_{\rm SN}\simeq 10^{51}~{}{\rm erg}$ into the surrounding gas. However, at high densities, the cooling radius is smaller than the resolution scale. In this case, injecting energy into the surrounding gas would result in the energy from the SN being artificially radiated away before it could accelerate gas in the snowplow phase. To compensate for this effect, we inject momentum in addition to thermal energy when the cooling radius is not resolved. This momentum feedback method is common practice in galaxy simulations (Hopkins et al., 2011, 2018; Kim & Ostriker, 2015; Kimm et al., 2015).
@@
$R_{\rm cool}\simeq$ $3.0\left(\frac{Z}{Z_{\odot}}\right)^{-0.082}\left(\frac{n_{\rm H}}{100~{}{\rm cm^{-2}}}\right)^{-0.42}~{}{\rm pc}$ (27)
$P_{\rm SN}\simeq$ $1420\left(\frac{Z}{Z_{\odot}}\right)^{-0.137}\left(\frac{n_{\rm H}}{100~{}{\rm cm^{-2}}}\right)^{-0.16}~{}M_{\odot}{\rm km/s}$ (28)
@@
Our approach is only an approximation of the full model recommended by Martizzi et al. (2015), which incorporates additional physical scales into the calculation of injected momentum and energy. Martizzi et al. (2015) also provide versions of Equations 27 and 28 for SN which occur in an inhomogeneous medium with $\mathcal{M}_{\rm turb}=30$. In an inhomogeneous medium, the cooling radius is larger by a factor of a few because the blast wave preferentially moves through low-density channels, but the terminal energy and momentum deposited in the ISM by the SN remnant do not change significantly. Ideally, one should use a model which depends continuously on the turbulent Mach number.
@@
Momentum feedback approximates the long-term effect of SNe on the surrounding ISM, but it fails to capture heating on short timescales in the unresolved Sedov-Taylor phase. Our results concerning SN feedback should be considered with caution, despite our physically-motivated subgrid prescription. We leave further discussion to Section 4.7.
@@ 2.5 Early feedback recipe
In addition to SN feedback, stellar feedback includes thermal pressure from photoionized gas, stellar winds, and radiation pressure. We call these early feedback processes because they begin immediately after a massive star forms, in contrast to SN feedback. Out of the early feedback processes, we only model thermal pressure from photoionized gas. We discuss the effects of other early feedback processes in Section 4.5.
In each cell containing a star cluster particle younger than $\tau_{\rm end}=20~{}{\rm Myr}$, we set the temperature $T_{\rm phot}=20000~{}{\rm K}$. This forces the gas into a fast cooling regime where it approaches the photoionization temperature within a few timesteps. In practice, we achieve this by injecting a thermal energy associated with a sound speed
@@
where we have assumed a $\gamma$ = 5/3 adiabatic EOS and mean molecular weight $\mu=0.6$.
Our early feedback recipe depends on the choice of $M_{\rm cl}$ because each star particle has the same effect on its parent cell, regardless of its mass. As the adopted value of $M_{\rm cl}$ decreases, less stellar mass is required to photoionize the parent cell, representing increasingly efficient photoionization feedback. We select a fiducial value of $M_{\rm cl}$ such that the H II region volume equals the volume of the most refined cell under typical star-forming conditions. In this case, a star particle is formed exactly when its H II region becomes resolved.
@@
Using data for main sequence stars given by Table 15.1 of Draine (2011), we fit $Q(m)$ to a simple analytic form for masses $\geq 15.6~{}M_{\odot}$ (App. B):
$Q(m)\simeq\left[8.37\times 10^{45}\left(\frac{m}{M_{\odot}}\right)^{2.12}-2.33\times 10^{48}\right]~{}{\rm cm^{-2}s^{-1}}$ (32)
Assuming a Chabrier IMF (Chabrier, 2003), we estimate $\expectationvalue{Q/m_{*}}\simeq 5.85\times 10^{46}~{}{\rm cm^{-2}s^{-1}}M_{\odot}^{-1}$. Higher mass stars contribute more ionizing photons, but have a lower abundance. In our calculation, most of the contribution to the ionizing photons comes from stars $\simeq 35~{}M_{\odot}$. Star clusters which are too small to sample the IMF up to this mass may have lower ionization rates.
@@
where $\alpha_{B}\simeq 2.58\times 10^{-13}~{}{\rm cm^{3}/s}$ is the case B recombination rate at an electron temperature $10^{4}~{}{\rm K}$ (Ferland et al., 1992). We use the on-the-spot approximation, assuming that electrons which recombine directly to the ground state do not contribute to the net ionization.
The stellar mass required to photoionize a cell is given by equating Strömgren sphere and cell volumes and solving for the star cluster mass, which yields a similar value to our fiducial star particle mass $4400~{}M_{\odot}$:
$\begin{split}M_{\rm cl}=\ &\frac{4\pi}{3}\frac{\alpha_{B}n_{\rm H}^{2}\Delta x_{\rm min}^{3}}{\expectationvalue{Q/m_{*}}}\\ \simeq\ &5000~{}M_{\odot}\left(\frac{n_{\rm H}}{100~{}{\rm cm^{-3}}}\right)^{2}\left(\frac{\Delta x}{10~{}{\rm pc}}\right)^{3}\left(\frac{\expectationvalue{Q/m_{*}}~{}{\rm cm^{2}s}M_{\odot}}{5.85\times 10^{46}}\right)^{-1}\end{split}$ (34)
In MDGs, stars form at significantly higher densities than the value $\sim 10^{2}~{}{\rm cm^{-2}}$ which appears in Equation 34. However, the strong turbulence in MDGs lowers the effective density relevant for photoionization because photoionizing photons preferentially escape through low-density channels in the gas. We can estimate the effective density using a PDF-weighted harmonic mean
@@
This is equivalent to computing the density associated with the average photon mean free path, which scales as $\propto\rho^{-1}$ for constant opacity. The result comports with our intuition that the effective density should decrease with increasing turbulent Mach number. This is similar in spirit to the more heuristic definition of the effective density in Appendix B of Faucher-Giguère et al. (2013). For the typical densities and turbulent Mach numbers at star-formation sites in our simulations ($n_{\rm H}\sim 10^{4}~{}{\rm cm^{-3}}$, $\mathcal{M}_{\rm turb}\sim 10$), the effective density is indeed $\sim 10^{2}~{}{\rm cm^{-3}}$.
Our crude model (single star particle mass) cannot account for the density dependence in Equation 34. We run several simulations with different star particle masses to determine how this parameter effects our results (Sec. 2.6). In future work, we will improve upon our early feedback model.
We also do not properly account for multiple star cluster particles in a cell. If one star cluster particle photoionizes a cell, then by the Strömgren sphere argument, multiple star particles should photoionize a region larger than a cell. In our simulations, these situations are instead handled by multiplying the injected thermal energy by the number of star cluster particles, artificially restricting the Strömgren sphere to the cell size.